OpenStudy (anonymous):

Can you please check my answer? Thanks. -Problem- Add or subtract. (2x+3/x-4) - (x-5/x+2) -My Answer- (x+8/-6) Do you agree?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

2x - x = 1x, 3 - (-5) = 8, -4 - 2 = -6

OpenStudy (mertsj):